# Matching Hadronization and Perturbative Evolution: The Cluster Model in Light of Infrared Shower Cutoff Dependence

### v1

@@ 3.2 Thrust Distribution, Factorization and $R$-Evolution
where $Q$ is the $e^{+}e^{-}$ c.m. energy, and the sum runs over all final state particles with momenta $\vec{p_{i}}$. The maximum defines the thrust axis $\vec{n_{t}}$. In the limit of vanishing hadron masses $\tau$ is identical to thrust. The difference is very small and does not play any role in the context of our studies. We therefore call $\tau$ also thrust sometimes in the rest of this article. Thrust is an IR safe and global shape variable and in the limit of small $\tau$, referred to as the dijet region, the events are characterized by two energetic back-to-back jets along the thrust axis. In this section we render factorization aspects of the coherent branching shower cutoff $Q_{0}$ and the resulting RG equation explicit for the thrust distribution in the dijet region. This RG equation serves as the basis of the concrete numerical studies we carry out in Secs. 7 for the default and our novel dynamical hadronization model.
@@
The point essential for our study is that the form of Eq. (4), which factorizes parton level and hadronization effects, is an umambiguous property of QCD independent of the scheme that is adopted to regulate IR momenta in the partonic distribution. In Ref. Hoang:2018zrp the NLL partonic thrust distribution was analyzed from the perspective of using the transverse momentum cut $Q_{0}$ that is employed in Herwig’s angular ordered parton shower to regulate IR momenta and keeping track of the dominant linear dependence on $Q_{0}$. [^7] In Ref. Hoang:2018zrp the analysis was carried out for the coherent branching algorithm as well in soft-collinear effective theory (SCET). The exact relation of NLL precision for the angular ordered parton shower and terms in the NLL$+{\cal O}(\alpha_{s})$ order counting in SCET was specified as well. It was found that only the large-angle soft radiation can cause a linear $Q_{0}$ sensitivity and that the partonic distribution $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q,Q_{0})$ in the presence of a finite value of $Q_{0}$ can be written as
@@
and where $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q)$ is the partonic thrust distribution for a vanishing IR cutoff shown in Eq. (4). Note that the simple form of Eq. (7) arises from a multipole expansion keeping the dominant linear dependence on $Q_{0}$ of the full result (which is also given in Ref. Hoang:2018zrp ). Upon convolution with the shape function this multipole expansion provides an excellent approximation to the full result. The strong coupling in the gap function $\Delta_{\rm soft}(Q_{0})$ is evaluated at the renormalization scale $\mu=Q_{0}$ since the gap function quantifies the effects of the unresolved (large-angle) soft radiation below the scale $Q_{0}$, which can only depend on the scale $Q_{0}$ in perturbation theory. The NLL precision of the coherent branching algorithm ensures that this $Q_{0}$ dependence is also realized by the Herwig simulation parton level results. The ${\cal O}(\alpha_{s}^{2})$ term indicated in Eq. (7) can only be specified once the cutoff prescription in the context of a more precise NNLL order shower evolution has been defined. Such a prescription is currently unknown and we therefore drop these higher order contributions from now on. The result for the gap function $\Delta_{\rm soft}(Q_{0})$ implies it satisfies the renormalization group equation
@@
This evolution equation describes a linear scale dependence and has been called $R$-evolution in Refs. Hoang:2008fs ; Hoang:2008yj ; Hoang:2009yr .888 The point of the $R$-evolution is that the perturbation series for $\Delta_{\rm soft}(R)$ contains an ${\cal O}(\Lambda_{\rm QCD})$ IR renormalon, but that Eqs. (8) and (9) are renormalon-free. The NLL partonic thrust distribution at two different shower cutoff values $Q_{0}$ and $Q_{0}^{\prime}$ are therefore related by the equality
@@ 3.3 Shower Cutoff Dependence for Thrust and Migration Matrix
At this point it is highly instructive to cross check the level of validity of relation (9) for the implementation of the angular-ordered parton shower in Herwig. As we already mentioned, Herwig’s angular-ordered parton shower is NLL precise for $e^{+}e^{-}$ event-shapes in the dijet limit. Even though the NLL precision is guaranteed conceptually, it is certainly useful to examine it for the practical implementation of the Herwig event generator. To this end, we consider the shower-cutoff dependence of the 2-jettiness cumulant
@@
Figure 1: Binned cumulant differences of the parton level thrust distribution defined in Eq. (16) generated by the Herwig 7.2 angular ordered parton shower for $Q=45$ GeV (left panel), $91.2$ GeV (middle panel) and $200$ GeV (right panel) and shower cutoff values $Q_{0}=1$ GeV (blue) $1.5$ GeV (green) and $1.75$ GeV (red) with respect to the reference cutoff $Q_{0,{\rm ref}}=1.25$ GeV.
@@
In Fig. 1 we have displayed the cumulant difference of Eq. (16) for $Q_{0}=1.0$ (blue), $1.25$ (orange), $1.5$ (green) and $1.75$ GeV (red) with respect to the reference result for $Q_{0}^{\prime}=1.25$ GeV for the hard scales $Q=45$, $91.2$ and $200$ GeV. We have employed bins of size $\Delta\hat{\tau}=0.02$, and each bin value is determined from the average of the cumulant at the respective upper and lower bin boundaries divided by the differential cross section of the bin. To visualize the impact of matching corrections which affect the 2-jettiness distribution outside the dijet region, four different Herwig matching settings have been used. The four different intensities for the same color correspond to (from darkest to the brightest color): (i) leading-order matrix element with matrix-element (ME) correction for hard QCD radiation (which is Herwig’s default); (ii) the same without ME correction; (iii) NLO matrix elements with multiplicative (POWHEG-type) matching and (iv) additive (MC@NLO-type) NLO matching, as available from the Matchbox framework Platzer:2011bc . The dotted horizontal lines correspond do $\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})$, where we use the strong coupling in the $\overline{\rm MS}$ scheme extracted from Herwig to solve the RG equation of Eq. (10). As expected, the impact of the matching and matrix element corrections increases for larger $\hat{\tau}$ values since the soft-collinear approximations in the coherent branching algorithms work more efficiently in the dijet region where $\hat{\tau}$ is small. We also see that the region in $\hat{\tau}$, where the matching corrections become sizeable, increases toward smaller $\hat{\tau}$ ranges for increasing c.m. energy $Q$. This indicates that the dijet 2-jettiness region around $\hat{\tau}=0$ decreases when the hard scattering scale gets larger and is consistent with the fact that the peak location and peak width scale with $\Lambda/Q$.
@@
We see that, except for the very first $\hat{\tau}$ bin, where the no-radiation events and the implementation dependent details of the shower-cutoff dependence are still being resolved (and therefore not relevant for testing Eq. (16)), the partonic cumulant differences obtained from Herwig’s angular ordered shower are nicely compatible with the QCD value of $\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})$ for $\hat{\tau}$ up to about $0.2$. The visible discrepancies are related to quadratic and higher order effects in $Q_{0}$ which the linear approximation used in Ref. Hoang:2018zrp and expressed in Eq. (6) does not capture. Overall, we can conclude that the factorization formula (4) should be applicable for 2-jettiness values $\hat{\tau}$ up to around $0.2$ which well includes the peak region and a large fraction of 2-jettiness distribution tail.
@@
The transfer matrix $T$ is determined by reading out, for each event, the thrust values at the true parton level ($\hat{\tau}$) and at the hadron level ($\tau$) and by determining the resulting 2-dimensional histrogram. Unitarity ensures that $\int\mathrm{d}\tau\,T(\tau,\hat{\tau},\{Q,Q_{0}\})=1$ for each parton level $\hat{\tau}$. Comparing to the factorization analogue in Eq. (11) we see the close relation between the shape function $QS_{\rm had}(Q(\tau-\hat{\tau}),Q_{0})$, which satisfies the analogous normalization condition, and the migration matrix $T(\tau,\hat{\tau},\{Q,Q_{0}\})$. Changing variables to $k=Q\tau$ and $\hat{k}=Q\hat{\tau}$ we define the rescaled MC parton-to-hadron level migration matrix function
@@ 5 Dynamical Cluster Hadronization Model
Within Herwig’s cluster hadronization model implementation the two essential features that influence to which extend the hadronization model dynamics can properly match the infrared features of the parton shower are the gluon mass $m_{g}$, the kinematics of the subsequent forced gluon splitting in the first step of the hadronization process, and the dynamics behind the cluster fission. For a fixed gluon mass value the problematic aspect is that the gluon splitting process carried out by the parton shower entails in contrast a nontrivial distribution of invariant masses of the $g\to q\bar{q}$ process. This distribution depends on the value of shower cut $Q_{0}$. But it is also clear that a fixed gluon mass value can only mimic some averaged features of the parton shower splitting, and this prohibits an exact matching to the parton shower. For the cluster fission process, where an additional light quark-anti-quark pair is produced from the vacuum, the parton shower analogue is a radiated gluon which afterwords branches into the light quark-antiquark pair. For the default cluster fission process, the dynamical aspects encoded in these parton shower processes are missing. It can therefore not be expected that fixing the parameters $m_{g}$, ${\rm{Cl}}_{\rm{max}}$, ${\rm{Cl}}_{\rm{pow}}$ and $\rm{PSplit}$ from the tuning procedure should result in an exact matching to the parton shower.
The motivation for the construction of the dynamical cluster hadronization model is to add these parton shower aspects back to the model implementation such that the ultraviolet aspects of the hadronization modelling have an improved compatibility to the infrared behavior of the parton shower. The essential novel aspects are (i) a dynamically generated distribution for the gluon mass $m_{g}$ and parton-shower-like kinematics for the forced splitting in the initial stage of the hadronization and (ii) the implementation of a parton-shower-like dynamics behind the cluster fission process. These are explained in the following two subsections.
@@ 5.2 Embedding Gluon Branching into Cluster Fission
The gluon with momentum $\tilde{g}^{\mu}$ now splits into a $q^{\prime}\bar{q}^{\prime}$ pair following the description of the forced gluon splitting in Sec. 5.1. However, the backwards direction $\bar{n}^{\mu}$ is given in Eq. (43). Furthermore, the analogue of the scale $\tilde{Q}_{g}$ (shown in Eq. (29)) which here we denote by $\tilde{Q}_{g}^{(f)}$ is related to the splitting variables $z$ and $\tilde{q}$ by the angular ordering relation111111 We note that there is also the option to use $\tilde{Q}_{g}^{(f)}$ as an additional tuning parameter that sets the scale of the gluon splitting in the fission independent of the (anti)quark branching. We have, however, not used this option in our analyses.
@@ 7.1 Shower Cutoff Independence of Hadron Level Observables
We have checked that these features are not only realized for 2-jettiness and event-shapes, on which most of our analytic insights concerning the NLL precision of Herwig’s angular ordered parton shower and of the partonic shower cut $Q_{0}$ dependence are based on, but actually for all jet and event-shape related observables that have been measured in $e^{+}e^{-}$ collisions in the past. In Figs. 16 and 17 shown in App. A this is demonstrated displaying the simulations results obtained for the default and the novel dynamical hadronization model, respectively, again for the shower cutoff values $Q_{0}=1$, $1.25$, $1.5$ and $1.75$ GeV for a number of selected observables for in comparison with actual LEP data gathered at $Q=44$, $Q=91$ and $Q=133$ GeV. We have displayed the results for a number of other event-shapes (thrust, heavy jet mass), jet resolution rates as well as charged particle multiplicities measured at the JADE, ALEPH, DELPHI and OPAL experiments MovillaFernandez:1997fr ; JADE:1999zar ; ALEPH:2003obs ; ALEPH:1996oqp ; ALEPH:1991ldi ; DELPHI:1996sen The respective Rivet analyses codes used for the generation of each of the histograms is quoted at the bottom of each panel. The lower sections of the individual panels show ratios with respect to the MC reference simulation results for $Q_{0,\rm{ref}}=1.25$ GeV. The improved shower cutoff $Q_{0}$-independence for the dynamical model in comparison to the default model for all observables is clearly visible. We emphasize again that we have checked that the results shown in Figs. 16 and 17 are representative for all available $e^{+}e^{-}$ Rivet analyses. We also stress that, overall, the quality of the description of the data provided by all the tunes we obtained for the novel dynamical hadronization model is the same as that of the standard Herwig release tune.

### v2

@@ 3.2 Thrust Distribution, Factorization and $R$-Evolution
where $Q$ is the $e^{+}e^{-}$ c.m. energy, and the sum runs over all final state particles with momenta $\vec{p_{i}}$. The maximum defines the thrust axis $\vec{n_{t}}$. In the limit of vanishing hadron masses $\tau$ is identical to thrust. The difference is very small and does not play any role in the context of our studies. We therefore call $\tau$ also thrust sometimes in the rest of this article. Thrust is an IR safe and global shape variable and in the limit of small $\tau$, referred to as the dijet region, the events are characterized by two energetic back-to-back jets along the thrust axis. In this section we render factorization aspects of the coherent branching shower cutoff $Q_{0}$ and the resulting RG equation explicit for the thrust distribution in the dijet region. This RG equation serves as the basis of the concrete numerical studies we carry out in Secs. 7 for the default and our novel dynamical hadronization model. As already indicated in the introduction, results along these lines are currently only available for jet mass related event-shapes such as 2-jettiness.
@@
The point essential for our study is that the form of Eq. (4), which factorizes parton level and hadronization effects, is an umambiguous property of QCD independent of the scheme that is adopted to regulate IR momenta in the partonic distribution. In Ref. Hoang:2018zrp the NLL partonic thrust distribution was analyzed from the perspective of using the transverse momentum cut $Q_{0}$ that is employed in Herwig’s angular ordered parton shower to regulate IR momenta and keeping track of the dominant linear dependence on $Q_{0}$. [^10] In Ref. Hoang:2018zrp the analysis was carried out for the coherent branching algorithm as well in soft-collinear effective theory (SCET). The exact relation of NLL precision for the angular ordered parton shower and terms in the NLL$+{\cal O}(\alpha_{s})$ order counting in SCET was specified as well. It was found that (a) to achieve NLL for the solution of the coherent branching algorithm only the splitting function $P_{qq}$ for the radiation of a gluon off a quark needs to be accounted for111111 This feature was already discussed a long time ago in Ref. Catani:1992ua . and that (b) only the large-angle soft radiation can cause a linear $Q_{0}$ sensitivity. The partonic distribution $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q,Q_{0})$ in the presence of a finite value of $Q_{0}$ can be written as
@@
and where $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q)$ is the partonic thrust distribution for a vanishing IR cutoff shown in Eq. (4). Note that the simple form of Eq. (7) arises from a multipole expansion keeping the dominant linear dependence on $Q_{0}$ of the full result (which is also given in Ref. Hoang:2018zrp ). Upon convolution with the shape function this multipole expansion provides an excellent approximation to the full result. The strong coupling in the gap function $\Delta_{\rm soft}(Q_{0})$ is evaluated at the renormalization scale $\mu=Q_{0}$ since the gap function quantifies the effects of the unresolved (large-angle) soft radiation below the scale $Q_{0}$, which can only depend on the scale $Q_{0}$ in perturbation theory. The NLL precision of the coherent branching algorithm ensures that this $Q_{0}$ dependence is also realized by the Herwig simulation parton level results.
The ${\cal O}(\alpha_{s}^{2})$ term indicated in Eq. (7) can only be specified once the cutoff prescription in the context of a more precise NNLL order shower evolution has been defined. Such a prescription is currently unknown and we therefore drop these higher order contributions from now on. The result for the gap function $\Delta_{\rm soft}(Q_{0})$ implies it satisfies the renormalization group equation
@@
This evolution equation describes a linear scale dependence and has been called $R$-evolution in Refs. Hoang:2008fs ; Hoang:2008yj ; Hoang:2009yr .121212 The point of the $R$-evolution is that the perturbation series for $\Delta_{\rm soft}(R)$ contains an ${\cal O}(\Lambda_{\rm QCD})$ IR renormalon, but that Eqs. (8) and (9) are renormalon-free. The NLL partonic thrust distribution at two different shower cutoff values $Q_{0}$ and $Q_{0}^{\prime}$ are therefore related by the equality
@@ 3.3 Shower Cutoff Dependence for Thrust and Migration Matrix
At this point it is highly instructive to cross check the level of validity of relation (9) for the implementation of the angular-ordered parton shower in Herwig. As we already mentioned, Herwig’s angular-ordered parton shower is NLL precise for $e^{+}e^{-}$ event-shapes in the dijet limit. Even though the NLL precision is guaranteed conceptually, it is certainly useful to examine it for the practical implementation of the Herwig event generator. Note that a similar analysis has already been carried out in Ref. Hoang:2018zrp . To this end, we consider the shower-cutoff dependence of the 2-jettiness cumulant
@@
Figure 1: Binned cumulant differences of the parton level thrust distribution defined in Eq. (16) generated by the Herwig 7.2 angular ordered parton shower for $Q=45$ GeV (left panel), $91.2$ GeV (middle panel) and $200$ GeV (right panel) and shower cutoff values $Q_{0}=1$ GeV (blue) $1.5$ GeV (green) and $1.75$ GeV (red) with respect to the reference cutoff $Q_{0,{\rm ref}}=1.25$ GeV. The dashed lines show the results expected from QCD factorization employing the $\overline{\rm MS}$ strong coupling extracted from Herwig.
@@
In Fig. 1 we have displayed the cumulant difference of Eq. (16) for $Q_{0}=1.0$ (blue), $1.25$ (orange), $1.5$ (green) and $1.75$ GeV (red) with respect to the reference result for $Q_{0}^{\prime}=1.25$ GeV for the hard scales $Q=45$, $91.2$ and $200$ GeV. We have employed bins of size $\Delta\hat{\tau}=0.02$, and each bin value is determined from the average of the cumulant at the respective upper and lower bin boundaries divided by the differential cross section of the bin. To visualize the impact of matching corrections which affect the 2-jettiness distribution outside the dijet region, four different Herwig matching settings have been used. The four different intensities for the same color correspond to (from darkest to the brightest color): (i) leading-order matrix element with matrix-element (ME) correction for hard QCD radiation (which is Herwig’s default); (ii) the same without ME correction; (iii) NLO matrix elements with multiplicative (POWHEG-type) matching and (iv) additive (MC@NLO-type) NLO matching, as available from the Matchbox framework Platzer:2011bc . The dotted horizontal lines correspond to $\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})$, where we use the strong coupling in the $\overline{\rm MS}$ scheme extracted from Herwig to solve the RG equation of Eq. (10). As expected, the impact of the matching and matrix element corrections increases for larger $\hat{\tau}$ values since the soft-collinear approximations in the coherent branching algorithms work more efficiently in the dijet region where $\hat{\tau}$ is small. We also see that the region in $\hat{\tau}$, where the matching corrections become sizeable, increases toward smaller $\hat{\tau}$ ranges for increasing c.m. energy $Q$. This indicates that the dijet 2-jettiness region around $\hat{\tau}=0$ decreases when the hard scattering scale gets larger and is consistent with the fact that the peak location and peak width scale with $\Lambda/Q$.
@@
We see that, except for the very first $\hat{\tau}$ bin, where the no-radiation events and the implementation dependent details of the shower-cutoff dependence are still being resolved (and therefore not relevant for testing Eq. (16)), the partonic cumulant differences obtained from Herwig’s angular ordered shower are nicely compatible with the QCD value of $\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})$ for $\hat{\tau}$ up to about $0.2$. The level of agreement is as expected from a comparison between the analytic result of Eq. (10) (which was obtained from an anomalous dimension at ${\cal O}(\alpha_{s})$ for the dominant linear cutoff scale effects in Ref. Hoang:2018zrp ) and the parton shower result (where higher power terms in $\alpha_{s}$ and $Q_{0}$ are generated as well). The observed deviations at the level of $10$ to $15\%$ are consistent with the expected size of the higher order NNLL QCD and quadratic $Q_{0}$ contributions generated in the parton shower simulation, but not contained in the analytic NLL QCD result for $\Delta_{\rm soft}$ in Eq. (10). We also refer to Ref.[2] for analytic solutions of the coherent branching algorithm for thrust with NLL precision. Overall, we can conclude that the factorization formula (4) should be applicable for 2-jettiness values $\hat{\tau}$ up to around $0.2$ which well includes the peak region and a large fraction of 2-jettiness distribution tail.
@@
The transfer matrix $T$ is determined by reading out, for each event, the thrust values at the true parton level ($\hat{\tau}$) and at the hadron level ($\tau$) and by determining the resulting 2-dimensional histogram. Unitarity ensures that $\int\mathrm{d}\tau\,T(\tau,\hat{\tau},\{Q,Q_{0}\})=1$ for each parton level $\hat{\tau}$. Comparing to the factorization analogue in Eq. (11) we see the close relation between the shape function $QS_{\rm had}(Q(\tau-\hat{\tau}),Q_{0})$, which satisfies the analogous normalization condition, and the migration matrix $T(\tau,\hat{\tau},\{Q,Q_{0}\})$. Changing variables to $k=Q\tau$ and $\hat{k}=Q\hat{\tau}$ we define the rescaled MC parton-to-hadron level migration matrix function
@@ 5 Dynamical Cluster Hadronization Model
Within Herwig’s cluster hadronization model implementation the two essential features that influence to which extend the hadronization model dynamics can properly match the infrared features of the parton shower are the gluon mass $m_{g}$, the kinematics of the subsequent forced gluon splitting in the first step of the hadronization process, and the dynamics behind the cluster fission. For a fixed gluon mass value the problematic aspect is that the gluon splitting process carried out by the parton shower entails in contrast a nontrivial distribution of invariant masses of the $g\to q\bar{q}$ process. This distribution depends on the value of shower cut $Q_{0}$. But it is also clear that a fixed gluon mass value can only mimic some averaged features of the parton shower splitting, and this prohibits an exact matching to the parton shower. For the cluster fission process, where an additional light quark-anti-quark pair is produced from the vacuum, the parton shower analogue is a radiated gluon which afterwards branches into the light quark-antiquark pair. For the default cluster fission process, the dynamical aspects encoded in these parton shower processes are missing. It can therefore not be expected that fixing the parameters $m_{g}$, ${\rm{Cl}}_{\rm{max}}$, ${\rm{Cl}}_{\rm{pow}}$ and $\rm{PSplit}$ from the tuning procedure should result in an exact matching to the parton shower.141414 Reference Chahal:2022rid also discusses a dynamical splitting of gluons depending on their colour connections to other perturbatively produced partons in the context of an implementation of a cluster hadronization model in the SHERPA event generator. From the algorithmic structure this model may be able to also relate the underlying splitting function to the shower evolution.
The motivation for the construction of the dynamical cluster hadronization model is to add these parton shower aspects back to the model implementation such that the ultraviolet aspects of the hadronization modelling have an improved compatibility to the infrared behavior of the parton shower. The essential novel aspects are (i) a dynamically generated distribution for the gluon mass $m_{g}$ and parton-shower-like kinematics for the forced splitting in the initial stage of the hadronization and (ii) the implementation of a parton-shower-like dynamics behind the cluster fission process. In particular, for (i) we supplement the splitting function $P_{g\to q\bar{q}}$ for the process $g\to q\bar{q}$ and base the cluster fission dynamics (ii) on the process $q\to qg\to qq^{\prime}\bar{q}^{\prime}$ governed by the splitting functions $P_{q\to qg}$ and $P_{g\to q\bar{q}}$. As the clusters are (currently) based only on quark-antiquark constituents, the model has no analogue to the gluon splitting $g\to gg$. These implementations are explained in the following two subsections.
@@ 5.2 Embedding Gluon Branching into Cluster Fission
The gluon with momentum $\tilde{g}^{\mu}$ now splits into a $q^{\prime}\bar{q}^{\prime}$ pair following the description of the forced gluon splitting in Sec. 5.1. However, the backwards direction $\bar{n}^{\mu}$ is given in Eq. (43). Furthermore, the analogue of the scale $\tilde{Q}_{g}$ (shown in Eq. (29)) which here we denote by $\tilde{Q}_{g}^{(f)}$ is related to the splitting variables $z$ and $\tilde{q}$ by the angular ordering relation161616 We note that there is also the option to use $\tilde{Q}_{g}^{(f)}$ as an additional tuning parameter that sets the scale of the gluon splitting in the fission independent of the (anti)quark branching. We have, however, not used this option in our analyses.
@@ 7.1 Shower Cutoff Independence of Hadron Level Observables
We have checked that these features are not only realized for 2-jettiness and event-shapes, on which most of our analytic insights concerning the NLL precision of Herwig’s angular ordered parton shower and of the partonic shower cut $Q_{0}$ dependence are based on, but actually for all jet and event-shape related observables that have been measured in $e^{+}e^{-}$ collisions in the past. In Figs. 16 and 17 shown in App. A this is demonstrated displaying the simulations results obtained for the default and the novel dynamical hadronization model, respectively, again for the shower cutoff values $Q_{0}=1$, $1.25$, $1.5$ and $1.75$ GeV for a number of selected observables for in comparison with actual LEP data gathered at $Q=44$, $Q=91$ and $Q=133$ GeV. We have displayed the results for a number of other event-shapes (thrust, heavy jet mass), jet resolution rates as well as charged particle multiplicities measured at the JADE, ALEPH, DELPHI and OPAL experiments MovillaFernandez:1997fr ; JADE:1999zar ; ALEPH:2003obs ; ALEPH:1996oqp ; ALEPH:1991ldi ; DELPHI:1996sen The respective Rivet analyses codes used for the generation of each of the histograms is quoted at the bottom of each panel. The lower sections of the individual panels show ratios with respect to the MC reference simulation results for $Q_{0,\rm{ref}}=1.25$ GeV. The improved shower cutoff $Q_{0}$-independence for the dynamical model in comparison to the default model for all observables is clearly visible. We emphasize again that we have checked that the results shown in Figs. 16 and 17 are representative for all available $e^{+}e^{-}$ Rivet analyses. We also stress that, overall, the quality of the description of the data provided by all the tunes we obtained for the novel dynamical hadronization model is the same as that of the standard Herwig release tune. [^20] The fact that this description is not the best description of the current data is a known feature of the angular ordered shower in the default, analytically controllable and logarithmically accurate setup we employ here. This has been partially addressed in Bewick:2019rbu ; Bewick:2021nhc and improvements are subject to ongoing work. We do not expect, however, that this will change any of our findings on the hadronization since the data has merely served as a reference for a realistic tuning campaign.